How do you solve this math problem 15x + 5y = 20

Found 2 solutions by addingup, Alan3354:
Answer by addingup(3677)

15x+5y= 20
15x= 20-5y
x= 20-5y/15 All numbers are divisible by 5, so let's simplify
x= 4-y/3

Answer by Alan3354(69443)

It's not a problem.
It's and equation.
------
Step 1 - know what you want to do.
Step 2 - say what you want to do.